# Robustness of Dynamic Quantum Control: Differential Sensitivity Bounds

### v1

@@ I Introduction
Indeed, the robust control of quantum systems in the presence of noise has been considered for a variety of technologically-relevant systems, among them quantum computational architectures based on trapped ions Valahu et al. (2022); spin Barnes et al. (2015) and superconducting qubits Baum et al. (2021); Werninghaus et al. (2021); atom-interferometric inertial sensors Saywell et al. (2023); Weidner et al. (2017); and linear quantum optics Xiang et al. (2020), among many others. There are a number of methods for direct determination and optimization of robust quantum controllers, among them composite pulses Wimperis (1994); Motzoi et al. (2009), pulse shaping Daems et al. (2013), geometric control Chen et al. (2023), minimization of the quantum Fisher information Poggi et al. (2023), hybrid gradient-descent/simulated annealing Ram et al. (2022), adapted Krotov for disorder-dressed evolution Araki et al. (2023), and algorithms based on Pareto optimization Bhutoria et al. (2022).
@@
Beyond generating potential controllers, analysis methods for determining the robustness of controllers post-optimization are important. Robustness measures enable systematic comparison of controllers and facilitate the selection of those most suitable. One such measure is the Robustness Infidelity Measure (RIM) Khalid et al. (2023a), the lowest order of which reduces to the average fidelity error of a controller estimated by sampling over perturbations, or through propagation of uncertainties Dalgaard et al. (2022). Robustness measures for quantum controllers have also been developed based on the control-theoretic measures of differential and logarithmic sensitivity O’Neil et al. (2023a); O’Neil et al. (2023a), but to date, these methods have been limited to static control, or so-called energy landscape control Schirmer et al. (2018a), where the control fields are time-independent and the optimal system evolves from a predetermined initial state to a desired final state. While static control is a promising novel paradigm for certain applications, it has limitations, and the majority of control schemes proposed to-date for quantum systems are based on dynamic control. Dynamic control is usually formulated in terms of piecewise-constant control amplitudes Khaneja et al. (2005a); Tannor et al. (1992), found via optimization with respect to a basis Caneva et al. (2011); Sørensen et al. (2018), or one of the robust controller optimization methods listed above; a survey of optimization methods can be found in Koch et al., 2022. The robustness of dynamic control designs is typically assessed via Monte-Carlo sampling, which can be computationally expensive Dalgaard et al. (2022); Valahu et al. (2022), especially for dynamic controls with similarly dynamic perturbations.
@@
In an effort to produce analytic methods for determining the robustness of dynamic controllers, the notion of differential sensitivity has recently been extended to dynamic control with time-dependent, piecewise-constant controls O’Neil et al. (2023b). The focus of this paper is on the application of these techniques to quantum gate optimization problems with the aim to establish performance guarantees and understand trade-offs (or lack thereof) between performance and robustness, focusing on a set of benchmark problems originally laid out in Ref. Machnes et al., 2011. In the years since the publication of the original problems, work has continued on the control of qubit registers for quantum device optimization Mortimer et al. (2021) and spintronic systems Marchukov et al. (2016); Iversen et al. (2020). In addition, spin chains Thompson et al. (2016) and general spin networks Gosset et al. (2015) have been proposed as a means of performing universal quantum computation. Nonetheless, the problems originally formulated as benchmark problems for optimization algorithms, serve as a varied and useful testbed for our methods. Although this work specifically considers gate implementation for closed quantum systems, the methods considered here can be extended to open quantum systems subject to decoherence and dissipation in a straightforward manner. Algorithms and methods to control such systems can be found in numerous works Koch (2016); Kallush et al. (2022); Bhutoria et al. (2022); Araki et al. (2023). Related recent work has also explored state transfer in spin chains coupled to quantum baths Nie et al. (2021).
@@
The main contribution of this work is to demonstrate the efficacy of differential sensitivity-based methods O’Neil et al. (2023b) as a tool to analyze the robustness of the quantum gate fidelity realized by piecewise-constant controls to parametric uncertainty. This technique provides a valuable, analytically-based post-selector for dynamic quantum controllers, complementing more common stochastic methods Khalid et al. (2023a); Koswara et al. (2021a). We also demonstrate how the system control time and time resolution of the controls affect both gate fidelity and robustness in linear qubit registers with Ising or Heisenberg-type coupling.
@@ II.1 System Dynamics
Within each interval $[t_{k},t_{k+1}]$, we restrict the $M$ control fields to take constant values. Then the control pulses are described as $f_{m}^{(k)}\in\mathbb{R}$, where $m$ denotes the control channel associated with $H_{m}$ and $k$ the time interval starting at $t_{k}$. The total Hamiltonian during time interval $[t_{k},t_{k+1}]$ is
@@
Let $\chi^{(k)}(t)$ be a constant pulse of unit magnitude that is only non-zero for the interval $[t_{k},t_{k+1}]$. Then, the total Hamiltonian can be written as the sum of $H^{(k)}$ over time
@@
For closed systems, the dynamics are governed by the time-dependent Schrödinger equation, written in the Heisenberg picture as
@@
is the solution to the Schrödinger equation on the interval $[t_{k},t_{k+1}]$, and $\Phi^{(\kappa,0)}$ is the concatenation of the $\kappa$ time-ordered state transition matrices. Without loss of generality, taking the initial unitary $U_{0}$ to be the identity $I_{N}$, the figure of merit is
@@ II.2 Range of Problems
Table 1: A summary of the gate fidelity problems considered. Problems $1$ to $4$, $7$, and $9$ are set to maintain notational consistency with Ref. Machnes et al., 2011. Problems $5$, $6$, and $8$ are newly selected problems formulated specifically for the robustness analysis presented here. Individual Qubit Addressability denotes a control architecture in which each qubit is subject to control by an independent external field. Simultaneous Control on All Qubits is the same external field applied to all qubits in the linear register. Initial Qubit Control Only refers to the application of an external field to only the first qubit in the register. See Section II.2 for details.
@@
As a testbed for our robustness analysis for piecewise-constant, dynamic controllers we build on the problems presented in Ref. Machnes et al., 2011. These problems provide a range of system sizes and control implementations for linear qubit registers with Ising or Heisenberg-type coupling. This makes the problem selection suitable for assessing the effect of different interaction types and control settings, such as local addressability versus global control, on the robustness of the controlled system while remaining within the subset of linear qubit register architectures with nearest-neighbor coupling.
We focus here on quantum gate implementation problems, ranging in complexity from two- to five-qubit gates. The full list of problems considered is summarized in Table 1. The gate operation times $t_{f}$ and number of time steps $\kappa$ were selected to facilitate synthesis of high-fidelity controllers $\left(\mathcal{F}(t_{f})>0.99\right)$. For most problems the values in Machnes et al., 2011 were chosen as a starting point for each problem, with gate operation times varied from the shortest times for which we can find controllers that achieve minimum performance criteria to larger $t_{f}$, where it is usually easier to find solutions. Similarly, for the number of time steps $\kappa$, we typically aim to start with values just large enough to be able to find controllers that meet certain minimum performance requirements and then increase $\kappa$, which increases the dimension of the search space and therefore (generally) the set of controllers that meet minimum performance requirements. For example, we may select minimum values for $t_{f}$ and $\kappa$ such that our optimization procedure yields controllers that achieve fidelities $\geq 0.99$ with at least some controllers achieving fidelities $>0.9999$.
@@
where $Q$ is the number of qubits, $\sigma_{\set{x,y,z}}^{(\ell)}$ denotes the $Q-$fold tensor product of $I_{2}$ with $\sigma_{\set{x,y,z}}$ in the $\ell$-th position, and $\{\alpha,\beta\}\in\{0,1\}$ differentiate the drift Hamiltonian for distinct problems as described below. The matrices $\sigma_{\set{x,y,z}}$ are the $2\times 2$ Pauli matrices, and $I_{2}$ is the identity. The second term in the drift Hamiltonian corresponds to local on-site potentials, while the first term represents couplings between adjacent qubits, which are assumed to be uniform and fixed. In the following, we choose energy in units of $\hbar J$ and time in units of $J^{-1}$, drop the factors of $\hbar J$. If we assume resonant driving fields, the local potentials can usually be eliminated by transforming to a rotating frame and making the rotating wave approximation, in which case the drift Hamiltonian simplifies to
@@
To assess whether the interaction type affects the sensitivity behavior we consider a qubit register with Heisenberg coupling, for which we set $\alpha=\beta=1$ in Eq. (9), again assuming full local control given by Eq. (10). Specifically, we compare the implementation of a three-qubit QFT gate for a Heisenberg-coupled system (Problem $5$) with its Ising equivalent (Problem $2$). To eliminate target gate dependence and amplify the control difficulty, a random unitary operator $U_{f}\in\mathbb{U}(8)$, with elements distributed by Haar measure on $\mathbb{U}(8)$ Mezzadri (2007), is chosen as a target gate for Problem $6$.
@@
Another aspect that could plausibly influence the robustness of the optimized controllers is the type of control. Individual spin addressability generally affords the most control, but this is not always required. For example, it often suffices to have control over a single qubit at one end of the register as the effect of the controls are propagated along the chain by Heisenberg (but not Ising-type) coupling Schirmer et al. (2008). To assess the effect of such restricted control, we include the implementation of a three-qubit QFT and random unitary gate for a Heisenberg-coupled register with control of the first qubit only (Problems $8$, $9$). We designate this control implementation as Initial Qubit Control Only in Table 1. For these problems, the control Hamiltonians reduce to $H_{1}=\tfrac{1}{2}\sigma_{x}^{(1)}$ and $H_{2}=\tfrac{1}{2}\sigma_{y}^{(1)}$. For Problem $8$ the target unitary is a QFT, while for Problem $9$ the target is the same random unitary used for Problem $6$.
@@
Selective addressing of qubits can be achieved in certain control settings such as laser control of trapped ions, atoms or NV centers, where a laser can be focused on individual qubits, or quantum registers where qubits are controlled by surface gate electrodes. However, direct selective addressing is often not possible. For example, in the microwave regime, focusing microwaves on individual qubits is challenging, and in molecular systems and applications involving nuclear magnetic resonance (NMR) or electron spin resonance (ESR), magnetic fields cannot be focused directly on a single nuclear or electron spin. In this case selectivity is typically achieved by frequency-selective addressing, either by taking advantage of existing chemical shifts or by applying electric or magnetic field gradients to ensure that different qubits have different resonance frequencies. However, frequency-selective addressing has drawbacks; in particular it requires that the control amplitudes must be modulated on time-scales that are slow compared to the frequency difference between qubits to minimize off-resonant excitation Nigmatullin and Schirmer (2009). In these cases it is often preferable to use a global control model. To cover this case, in Problem $7$ we consider a linear qubit register with Ising coupling between adjacent qubits as before but adding a position-dependent Stark or Zeeman shift, which leads to a drift Hamiltonian where the local onsite potentials do not vanish:
@@ II.3 Controller Synthesis
The focus of this paper is the analysis of the robustness of controllers. The techniques presented are independent of the synthesis method used and can be applied to controllers obtained from arbitrary algorithms. Here, we use standard algorithms similar to those used in Ref. Floether et al., 2012 to generate the controllers. Unless otherwise stated, the controllers analyzed in this paper were generated by unconstrained optimization in MATLAB using the fminunc function, with the goal of producing an optimal sequence of control fields $\set{f_{m}^{(k)}}_{k=0}^{\kappa-1}$ that minimize $\varepsilon$ with $\kappa$ time steps. The initial condition for the optimization problem was an $M\times\kappa$ array of initial control fields for each control channel and time interval $[t_{k},t_{k+1}]$. The initial values were either drawn randomly from a uniform distribution or a standard normal distribution, and for most problems one initial condition with all fields set to 00 was also included.
@@ III Robust Performance in the Presence of Uncertainty
Predictions of any model of an experimental system are subject to uncertainty in practice. Understanding the sensitivity of the control performance with regard to various uncertainties is of utmost importance. This can be done numerically by Monte Carlo simulations, randomly varying model parameters and performing statistical analysis. However, this is computationally intensive and may not provide deeper insight into which uncertainties are most critical in terms of robustness. Here, we consider an alternative approach that quantifies the effect of structured perturbations to the Hamiltonian on the fidelity error in terms of the differential sensitivity of the fidelity error O’Neil et al. (2023b). We establish the uncertainty model for the sensitivity analysis, present the differential sensitivity of the fidelity error for dynamic controls, and discuss upper bounds on the latter that are agnostic to the exact structure of the uncertainty.
@@ III.1 Uncertainty Model
Uncertainties in model parameters often lead to structured perturbations to the Hamiltonian. For example, uncertainty in the $J$-coupling between spins will manifest itself in a structured uncertainty to the drift Hamiltonian. The structure depends on which couplings are affected, if they are affected independently or collectively, etc. Formally, we model structured uncertainties to the Hamiltonian $H(t)$ as $\delta\hat{H}_{\mu}$, where $\delta\in[\delta_{1},\delta_{2}]$ is the scalar deviation of the uncertain parameter from its nominal value, and $\hat{H}_{\mu}$ is a Hermitian matrix, normalized with respect to the Frobenius norm such that $\|\hat{H}_{\mu}\|_{F}=1$. For piecewise-constant controls, we write the perturbed Hamiltonian $\tilde{H}(t)$ as a sum of the perturbed Hamiltonians $\tilde{H}^{(k)}$ for the $k$th time step:
@@
$\tilde{H}^{(k)}$ $=H_{0}+\sum_{m=1}^{M}H_{m}f_{m}^{(k)}+\delta\hat{H}_{\mu}\alpha_{\mu}^{(k)}$ (14b)
where $\alpha_{\mu}^{(k)}$ scales the perturbation as described below. This formulation is very general. For example, we could model an uncertainty in a single $J$-coupling between qubits $1$ and $2$, resulting in a deviation from the nominal value $J$, by setting $\hat{H}_{\mu}=\frac{1}{2}\sigma_{z}^{(1)}\sigma_{z}^{(2)}$ with $\delta=J_{12}-J$. We focus on three main cases here: (1) uncertainty in the drift Hamiltonian $H_{0}$, (2) uncertainty in a control Hamiltonian $H_{m}$, and (3) linear combinations of the drift and control Hamiltonian uncertainty.
In case (1) we set $\tilde{H}_{0}=H_{0}+\delta\hat{H}_{0}$, so $\hat{H}_{\mu}=\hat{H}_{0}$ and $\alpha_{\mu}^{(k)}=1$ for all $k$, where $\hat{H}_{0}=H_{0}/\|H_{0}\|_{F}$ is the normalized structure matrix for $H_{0}$. We choose to consider a collective uncertainty in the drift Hamiltonian to reflect uncertainty due to modeling error in the case uniform coupling. This analysis can be extended to consider a subset of uncertain couplings, as mentioned above.
In case (2) we set $\tilde{H}_{m}f_{m}^{(k)}=(H_{m}+\delta\hat{H}_{m})f_{m}^{(k)}$ so that the final term in Eq. (III.1) is $\delta\hat{H}_{m}f_{m}^{(k)}$, where the normalized structure matrix for $H_{m}$ is $\hat{H}_{\mu}=\hat{H}_{m}$, and $\alpha_{\mu}^{(k)}=f_{m}^{(k)}$ with $\mu=m$.
In case (3) a combination of structured perturbations with uncertain Hamiltonian at time step $k$
$\tilde{H}^{(k)}=H^{(k)}+\delta\sum_{m=0}^{M}s_{m}\hat{H}_{m}\alpha_{m}^{(k)},$ (15)
is considered, where $s_{m}$ are scalar weights and $\mathbf{s}_{\mu}=(s_{0},\ldots,s_{M})$ must be a unit vector.

### v2

@@ I Introduction
Indeed, the robust control of quantum systems in the presence of noise has been considered for a variety of technologically-relevant systems, among them quantum computational architectures based on trapped ions [1]; spin [2] and superconducting qubits [3, 4]; atom-interferometric inertial sensors [5, 6]; and linear quantum optics [7], among many others. There are a number of methods for direct determination and optimization of robust quantum controllers, among them composite pulses [8, 9], pulse shaping [10], geometric control [11], minimization of the quantum Fisher information [12], hybrid gradient-descent/simulated annealing [13], adapted Krotov for disorder-dressed evolution [14], and algorithms based on Pareto optimization [15].
@@
Beyond generating potential controllers, analysis methods for determining the robustness of controllers post-optimization are important. Robustness measures enable systematic comparison of controllers and facilitate the selection of those most suitable. One such measure is the Robustness Infidelity Measure (RIM) [16], the lowest order of which reduces to the average fidelity error of a controller estimated by sampling over perturbations, or through propagation of uncertainties [17]. Robustness measures for quantum controllers have also been developed based on the control-theoretic measures of differential and logarithmic sensitivity [18, 19], but to date, these methods have been limited to static control, or so-called energy landscape control [20], where the control fields are time-independent and the optimal system evolves from a predetermined initial state to a desired final state. While static control is a promising novel paradigm for certain applications, it has limitations, and the majority of control schemes proposed to-date for quantum systems are based on dynamic control. Dynamic control is usually formulated in terms of piecewise-constant control amplitudes [21, 22], found via optimization with respect to a basis [23, 24], or one of the robust controller optimization methods listed above; a survey of optimization methods can be found in 25. The robustness of dynamic control designs is typically assessed via Monte-Carlo sampling, which can be computationally expensive [17, 1], especially for dynamic controls with similarly dynamic perturbations.
@@
In an effort to produce analytic methods for determining the robustness of dynamic controllers, the notion of differential sensitivity has recently been extended to dynamic control with time-dependent, piecewise-constant controls [26]. The focus of this paper is on the application of these techniques to quantum gate optimization problems with the aim to establish performance guarantees and understand trade-offs (or lack thereof) between performance and robustness, focusing on a set of benchmark problems originally laid out in Ref. 27. In the years since the publication of the original problems, work has continued on the control of qubit registers for quantum device optimization [28] and spintronic systems [29, 30]. In addition, spin chains [31] and general spin networks [32] have been proposed as a means of performing universal quantum computation. Nonetheless, the problems originally formulated as benchmark problems for optimization algorithms, serve as a varied and useful testbed for our methods. Although this work specifically considers gate implementation for closed quantum systems, the methods considered here can be extended to open quantum systems subject to decoherence and dissipation in a straightforward manner. Algorithms and methods to control such systems can be found in numerous works [33, 34, 15, 14]. Related recent work has also explored state transfer in spin chains coupled to quantum baths [35].
@@
The main contribution of this work is to demonstrate the efficacy of differential sensitivity-based methods [26] as a tool to analyze the robustness of the quantum gate fidelity realized by piecewise-constant controls to parametric uncertainty. This technique provides a valuable, analytically-based post-selector for dynamic quantum controllers, complementing more common stochastic methods [16, 36]. We also demonstrate how the system control time and time resolution of the controls affect both gate fidelity and robustness in linear qubit registers with Ising or Heisenberg-type coupling.
@@ II.1 System Dynamics
Within each interval $[t_{k},t_{k+1})$, we restrict the $M$ control fields to take constant values. Then the control pulses are described as $f_{m}^{(k)}\in\mathbb{R}$, where $m$ denotes the control channel associated with $H_{m}$ and $k$ the time interval starting at $t_{k}$. The total Hamiltonian during time interval $[t_{k},t_{k+1})$ is
@@
Let $\chi^{(k)}(t)$ be a constant pulse of unit magnitude that is only non-zero for the interval $[t_{k},t_{k+1})$. Then, the total Hamiltonian can be written as the sum of $H^{(k)}$ over time
@@
For closed systems, the dynamics are governed by the time-dependent Schrödinger equation
@@
is the solution to the Schrödinger equation on the interval $[t_{k},t_{k+1})$, and $\Phi^{(\kappa,0)}$ is the concatenation of the $\kappa$ time-ordered state transition matrices. Without loss of generality, taking the initial unitary $U_{0}$ to be the identity $I_{N}$, the figure of merit is
@@ II.2 Range of Problems
Table 1: A summary of the gate fidelity problems considered. Problems $1$ to $4$, $7$, and $9$ are set to maintain notational consistency with Ref. 27. Problems $5$, $6$, and $8$ are newly selected problems formulated specifically for the robustness analysis presented here. Individual Qubit Addressability denotes a control architecture in which each qubit is subject to control by an independent external field. Simultaneous Control on All Qubits is the same external field applied to all qubits in the linear register. Initial Qubit Control Only refers to the application of an external field to only the first qubit in the register. See Section II.2 for details.
@@
As a testbed for our robustness analysis for piecewise-constant, dynamic controllers we build on the problems presented in Ref. 27. These problems provide a range of system sizes and control implementations for linear qubit registers with Ising or Heisenberg-type coupling. This makes the problem selection suitable for assessing the effect of different interaction types and control settings, such as local addressability versus global control, on the robustness of the controlled system while remaining within the subset of linear qubit register architectures with nearest-neighbor coupling.
We focus here on quantum gate implementation problems, ranging in complexity from two- to five-qubit gates. The full list of problems considered is summarized in Table 1. The gate operation times $t_{f}$ and number of time steps $\kappa$ were selected to facilitate synthesis of high-fidelity controllers $\left(\mathcal{F}(t_{f})>0.99\right)$. For most problems the values in 27 were chosen as a starting point for each problem, with gate operation times varied from the shortest times for which we can find controllers that achieve minimum performance criteria to larger $t_{f}$, where it is usually easier to find solutions. Similarly, for the number of time steps $\kappa$, we typically aim to start with values just large enough to be able to find controllers that meet certain minimum performance requirements and then increase $\kappa$, which increases the dimension of the search space and therefore (generally) the set of controllers that meet minimum performance requirements. For example, we may select minimum values for $t_{f}$ and $\kappa$ such that our optimization procedure yields controllers that achieve fidelities $\geq 0.99$ with at least some controllers achieving fidelities $>0.9999$.
@@
where $Q$ is the number of qubits, $\sigma_{\set{x,y,z}}^{(\ell)}$ denotes the $Q-$fold tensor product of $I_{2}$ with $\sigma_{\set{x,y,z}}$ in the $\ell$-th position, and $\{\alpha,\beta\}\in\{0,1\}$ differentiate the drift Hamiltonian for distinct problems as described below. The matrices $\sigma_{\set{x,y,z}}$ are the $2\times 2$ Pauli matrices, and $I_{2}$ is the identity. The second term in the drift Hamiltonian corresponds to local on-site potentials with transition frequencies $\omega_{\ell}$, while the first term represents couplings between adjacent qubits, which are assumed to be uniform and fixed. In the following, we choose energy in units of $\hbar J$ and time in units of $J^{-1}$, drop the factors of $\hbar J$. If we assume that the transition frequencies are sufficiently well-separated to prevent off-resonant excitation and the control pulses vary much more slowly than the transition frequencies, then we can transform into a rotating frame and employ the rotating wave approximation in which the counter-rotating terms cancel over the cycle of a control pulse [37] to eliminate the local potentials, in which case the drift Hamiltonian simplifies to
@@
To assess whether the interaction type affects the sensitivity behavior we consider a qubit register with Heisenberg coupling, for which we set $\alpha=\beta=1$ in Eq. (9), again assuming full local control given by Eq. (10). Specifically, we compare the implementation of a three-qubit QFT gate for a Heisenberg-coupled system (Problem $5$) with its Ising equivalent (Problem $2$). To eliminate target gate dependence and amplify the control difficulty, a random unitary operator $U_{f}\in\mathbb{U}(8)$, with elements distributed by Haar measure on $\mathbb{U}(8)$ [38, 39, 40], is chosen as a target gate for Problem $6$.
@@
Another aspect that could plausibly influence the robustness of the optimized controllers is the type of control. Individual spin addressability generally affords the most control, but this is not always required. For example, it often suffices to have control over a single qubit at one end of the register as the effect of the controls are propagated along the chain by Heisenberg (but not Ising-type) coupling [41]. To assess the effect of such restricted control, we include the implementation of a three-qubit QFT and random unitary gate for a Heisenberg-coupled register with control of the first qubit only (Problems $8$, $9$). We designate this control implementation as Initial Qubit Control Only in Table 1. For these problems, the control Hamiltonians reduce to $H_{1}=\tfrac{1}{2}\sigma_{x}^{(1)}$ and $H_{2}=\tfrac{1}{2}\sigma_{y}^{(1)}$. For Problem $8$ the target unitary is a QFT, while for Problem $9$ the target is the same random unitary used for Problem $6$.
@@
Selective addressing of qubits can be achieved in certain control settings such as laser control of trapped ions, atoms or NV centers, where a laser can be focused on individual qubits, or quantum registers where qubits are controlled by surface gate electrodes. However, direct selective addressing is often not possible. For example, in the microwave regime, focusing microwaves on individual qubits is challenging, and in molecular systems and applications involving nuclear magnetic resonance (NMR) or electron spin resonance (ESR), magnetic fields cannot be focused directly on a single nuclear or electron spin. In this case selectivity is typically achieved by frequency-selective addressing, either by taking advantage of existing chemical shifts or by applying electric or magnetic field gradients to ensure that different qubits have different resonance frequencies. However, frequency-selective addressing has drawbacks; in particular it requires that the control amplitudes must be modulated on time-scales that are slow compared to the frequency difference between qubits to minimize off-resonant excitation [42]. In these cases it is often preferable to use a global control model. To cover this case, in Problem $7$ we consider a linear qubit register with Ising coupling between adjacent qubits as before but adding a position-dependent Stark or Zeeman shift, which leads to a drift Hamiltonian where the local onsite potentials do not vanish:
@@ II.3 Controller Synthesis
The focus of this paper is the analysis of the robustness of controllers. The techniques presented are independent of the synthesis method used and can be applied to controllers obtained from arbitrary algorithms. Here, we use standard algorithms similar to those used in Ref. 43 to generate the controllers. Unless otherwise stated, the controllers analyzed in this paper were generated by unconstrained optimization in MATLAB using the fminunc function, with the goal of producing an optimal sequence of control fields $\set{f_{m}^{(k)}}_{k=0}^{\kappa-1}$ that minimize $\varepsilon$ with $\kappa$ time steps. The initial condition for the optimization problem was an $M\times\kappa$ array of initial control fields for each control channel and time interval $[t_{k},t_{k+1})$. The initial values were either drawn randomly from a uniform distribution or a standard normal distribution, and for most problems one initial condition with all fields set to 00 was also included.
@@ III Robust Performance in the Presence of Uncertainty
Predictions of any model of an experimental system are subject to uncertainty in practice. Understanding the sensitivity of the control performance with regard to various uncertainties is of utmost importance. This can be done numerically by Monte Carlo simulations, randomly varying model parameters and performing statistical analysis. However, this is computationally intensive and may not provide deeper insight into which uncertainties are most critical in terms of robustness. Here, we consider an alternative approach that quantifies the effect of structured perturbations to the Hamiltonian on the fidelity error in terms of the differential sensitivity of the fidelity error [26]. We establish the uncertainty model for the sensitivity analysis, present the differential sensitivity of the fidelity error for dynamic controls, and discuss upper bounds on the latter that are agnostic to the exact structure of the uncertainty.
@@ III.1 Uncertainty Model
Uncertainties in model parameters often lead to structured perturbations to the Hamiltonian. For example, uncertainty in the $J$-coupling between spins will manifest itself in a structured uncertainty to the drift Hamiltonian. The structure depends on which couplings are affected, if they are affected independently or collectively, etc. Formally, we model structured uncertainties to the Hamiltonian $H(t)$ in the $k$th time-step as $\delta H^{(k)}_{\mu}$, where $\delta\in[\delta_{1},\delta_{2}]$ is the scalar deviation of the uncertain (set of) parameter(s) from their nominal values, and
$H^{(k)}_{\mu}:=\sum_{m=0}^{M}s_{m}^{(k)}\hat{H}_{m}\alpha_{m}^{(k)}$
represents the scaled structure of the uncertainty in time step $k$. Each $\hat{H}_{m}$ is a Hermitian matrix, normalized with respect to the Frobenius norm such that $\|\hat{H}_{m}\|_{F}=1$. Specifically, $\hat{H}_{0}$ models the uncertainty in the drift Hamiltonian, while the $\hat{H}_{m}$ for $1\leq m\leq M$ model the uncertainty in the control Hamiltonians. The set $\set{s^{(k)}_{m}}$ are scalar weights, normalized so that $\mathbf{s}^{(k)}_{\mu}=(s^{(k)}_{0},\ldots,s^{(k)}_{M})$ retains normalization of the un-scaled structure $\sum_{m=0}^{M}s^{(k)}_{m}\hat{H}_{m}$. The scalars $\alpha_{m}^{(k)}$ scale the structure based on the uncertainty type considered and described below.
We write the perturbed Hamiltonian $\tilde{H}(t)$ as a sum of the perturbed Hamiltonians $\tilde{H}^{(k)}$ for the $k$th time step:
@@
$\tilde{H}^{(k)}$ $=H_{0}+\sum_{m=1}^{M}H_{m}f_{m}^{(k)}+\delta{H}_{\mu}^{(k)}$ (14b)
$=H^{(k)}+\delta\sum_{m=0}^{M}s^{(k)}_{m}\hat{H}_{m}\alpha_{m}^{(k)}.$ (14c)
For uncertainty in the drift Hamiltonian, we have $\tilde{H}_{0}=H_{0}+\delta s^{(k)}_{0}\hat{H}_{0}$ so that $\alpha_{0}^{(k)}=1$ for all $k$. For uncertainty in the $m$th control Hamiltonian over time step $k$ we have $\tilde{H}_{m}f_{m}^{(k)}=(H_{m}+\delta s^{(k)}_{m}\hat{H}_{m})f_{m}^{(k)}=H_{m}f_{m}^{(k)}+\delta s^{(k)}_{m}\hat{H}_{m}f^{(k)}_{m}$, so that $\alpha_{m}^{(k)}=f_{m}^{(k)}$.
This formulation is very general. For example, we could model an uncertainty in a single $J$-coupling between qubits $1$ and $2$, resulting in a deviation from the nominal value $J$, by setting ${H}_{\mu}=\frac{1}{2}\sigma_{z}^{(1)}\sigma_{z}^{(2)}$ with $\delta=J_{12}-J$.